A circular iron ring of mean diameter 25 cm and cross-sectional area 9 π‘π‘š2 is wound with a coil of 100 turns and carries a current of 1.5 A. The relative permeability of iron is 2000. Calculate the amount of flux produced in the ring


1
Expert's answer
2022-03-10T18:01:29-0500

"B=\frac{\mu \mu_0 N I}{\pi d},"

"\Phi=SB=\frac{\mu \mu_0 NI S}{\pi d}=432~\mu Wb."